### v2

# GenRecal: Generation after Recalibration from Large to Small Vision-Language Models

Abstract
Recent advancements in vision-language models (VLMs) have leveraged large language models (LLMs) to achieve performance on par with closed-source systems like GPT-4V. However, deploying these models in real-world scenarios, particularly on resource-constrained devices, remains challenging due to their substantial computational demands. This has spurred interest in distilling knowledge from large VLMs into smaller, more efficient counterparts. A key challenge arises here from the diversity of VLM architectures, which are built on different LLMs and employ varying token types—differing in vocabulary size, token splits, and token index ordering. To address this challenge of limitation to a specific VLM type, we present Generation after Recalibration (GenRecal), a general-purpose distillation framework for VLMs. GenRecal incorporates a Recalibrator that aligns and adapts feature representations between heterogeneous VLMs, enabling effective knowledge transfer across different types of VLMs. Through extensive experiments on multiple challenging benchmarks, we demonstrate that GenRecal significantly improves baseline performances, eventually outperforming large-scale open- and closed-source VLMs.
Figure 1: (Left) Visualizing the token indices of a given image and text prompt and representing the possibility of distillation among various VLM pair combinations, comparing traditional distillation with our proposed distillation framework, GenRecal. Note that the parentheses mean each VLM’s LLM tokenizer, ‘…’ indicates the placement of image features, and the number of these features varies depending on the image embedding strategy. (Right) Comparing the performance of a challenging evaluation benchmark, $M$-Vet [yu2023mm], with [A] baseline, [B] SFT on the baseline, [C] traditional distillation and [D] GenRecal from same token types of large VLMs, and GenRecal with more powerful [E] large and [F] small VLMs.
00footnotetext: * Work Done during Internship.
Figure 2: (Left) Comparison of the challenging benchmark performances, MMB [liu2023mmbench], $M$-Vet [yu2023mm], MMMU [yue2023mmmu], and MMMU-Pro [yue2024mmmu] by changing large VLMs. The more powerful large VLMs we select, the greater the performance improvement we can achieve. (Right) Comparing the performance of the challenging benchmark: MMMU [yue2023mmmu], with GenRecal and various VLMs across model sizes. Note that all the experiments in Fig. 1 and Fig. 2 are conducted on the equal training dataset.
1 Introduction
Vision-language models (VLMs) have emerged as powerful tools for understanding and processing multimodal information, enabling tasks such as image captioning and visual question answering [tan2019lxmert, blip]. Recent advancements in VLMs [alayrac2022flamingo, blip2, liu2023improved] have leveraged large-scale language models (LLMs) to enhance reasoning and generation capabilities by incorporating extensive textual knowledge [2019t5, soldaini-etal-2024-dolma]. In pursuit of superior performance, state-of-the-art VLMs [li2024llava, wang2024qwen2vl, deitke2024molmo, nvlm2024, chen2024expanding] now integrate scaled-up LLMs such as 72B model parameters, achieving results comparable to proprietary models like GPT-4V [gptsyscard] and Claude-3.5 Sonnet [claude3series2024].
However, the increasing scale of recent VLMs introduces substantial computational overhead, limiting their practicality in real-world scenarios—particularly for on-device deployment. To mitigate this challenge, recent work has explored distillation techniques [cai2024llava, shu2024llava, feng2024align] that transfer knowledge from large (teacher) VLMs to smaller (student) ones. Yet, existing distillation methods suffer from a fundamental limitation: they typically assume that teacher and student produce sequences of equal token length, enabling token-level distance metrics such as KL divergence. This assumption breaks down when the two models use different tokenizers or adopt different image-splitting mechanisms (see Appendix A), making the distillation process infeasible. Consequently, the range of viable teacher–student VLM pairs becomes severely restricted (see Appendix B). In other words, current distillation approaches only operate when the teacher and student share identical vocabulary sizes, token split, and token index ordering schemes (referred to as token types), as illustrated on the left side of Fig. 1. These differences naturally lead to mismatched input or output token lengths across VLMs, preventing effective distillation. Note that even VLMs within the same family may employ different token types, further hindering compatibility.
Figure 3: Overview of the GenRecal architecture and its training stages. We denote the question and answer tokens from the small VLM as $q_{s}$ and $a_{s}$, and those from the large VLM as $q_{l}$ and $a_{l}$. For simplicity, the vision encoder and image-token projector are omitted from the illustration. Note that Recalibrator is used only during training as a bridge between heterogeneous VLMs. During inference, Recalibrator is removed, leaving the small VLM’s architecture unchanged and adding no extra computational cost.
To overcome this limitation, we propose Generation after Recalibration (GenRecal), a general-purpose distillation framework for token types-agnostic VLM distillation. At its core, GenRecal introduces a Recalibrator that aligns and adapts the feature representations of small VLMs with those of large VLMs. This design is inspired by prior studies demonstrating that word embeddings from different models can be linearly mapped into a shared embedding space [mikolov2013exploiting, smith2017offline, conneau2017word]. Unlike approaches that align only word embeddings, Recalibrator performs joint visual–linguistic alignment on the hidden representations before the language head. By projecting small VLM features into large VLMs’ feature space as shared representation, Recalibrator effectively bridges the gap between the large and small VLMs, enabling general-purpose knowledge transfer.
To demonstrate the necessity of GenRecal, we first compare its performance with that of traditional distillation, as shown in experiment types C and D of Fig. 1. Although both experiments employ the same token types of VLMs—Qwen2-VL-72B as the teacher and Qwen2-VL-7B as the student—GenRecal consistently outperforms the traditional distillation method implemented by LLaVA-KD [cai2024llava]. This result highlights that aligning feature representations between large and small VLMs is crucial for minimizing information loss during distillation, even when the models are homogeneous. Moreover, when we replace the large VLM from Qwen2-VL-72B with the more powerful InternVL2.5-78B, GenRecal achieves enhanced performance. Subsequently, replacing the small VLM with InternVL2.5-8B leads to an even stronger model, as summarized on the right side of Fig. 1. These experiments highlight the significance of GenRecal because traditional methods cannot handle VLM pairs with different token types.
Our contributions can be summarized as follows:
  • Token Types-agnostic Recalibration: GenRecal employs Recalibrator to align and adapt feature representations of large and small VLMs, enabling general-purpose distillation across various token types of vocabulary size, token splits, and token index ordering.
  • Broad Applicability: GenRecal is compatible with a wide range of VLM architectures across different model sizes, overcoming the challenge of selecting large VLMs that have different token types and demonstrating its practicality for real-world deployment in resource-constrained settings.
2 Related Work
Large-scale VLMs such as NVLM-72B [nvlm2024], Qwen2-VL-72B [wang2024qwen2vl], and InternVL2.5-78B [chen2024expanding] have closely reached the performances of GPT-4V [gptsyscard] and Claude-3.5 Sonnet [claude3series2024] since the evolution of VLMs (see Appendix C). However, they are imposing a significant computational burden in real-world applications, such as on-device processing. Hence, it is necessary to develop small VLMs to be embedded on the light devices, and this demand has led to the emergence of two types of distillation approaches. First one is constructing visual instruction tuning datasets [li2024llava, gu2024infinity, zhang2024beyond, li2024cvlm, chen2023can, hu2024mplug, wang2024enhancing, Sujet-Finance-QA-Vision-100k] from large-scale VLMs to train small VLMs with those curated datasets. Second one is conducting distillation: LLaVA-MoD [shu2024llava] conducts logits-based distillation on a mixture of expert (MoE) [shazeer2017, riquelme2021scaling] architectures. LLaVA-KD [cai2024llava] also uses logits-based distillation but presents three-stage training process where trainable parameters are different for each stage to effectively warm-up and distill the knowledge. Align-KD [feng2024align] distills both vision encoder features and decoder logits, while encouraging the student VLMs’ first decoder layer’s attention map to resemble that of the large VLM. MoVE-KD [cao2025move] employs multiple vision encoders to large VLMs and utilize a mixture of LoRA [hu2021lora] to small VLM. It distills visual information of large VLMs by using a visual attention map. Besides, there are earlier distillation studies (see Appendix C), but they do not consider either token-derived or different token type distillation.
[…]
Cambrian-1-8B [tong2024cambrian] 73.0 73.3 49.0 75.9 - 48.0 42.7 - 44.9 59.7 60.0
Cambrian-1-13B [tong2024cambrian] 73.6 73.8 48.0 75.7 - 48.9 40.0 - 43.1 60.0 58.6
Eagle-8B [shi2024eagle] 76.1 80.1 52.7 75.9 - 33.3 43.8 - 22.4 57.5 63.8
Eagle-13B [shi2024eagle] 74.0 77.6 54.4 75.7 - 42.6 41.6 - 21.8 60.2 62.9
VILA1.5-8B [lin2023vila] 58.8 - 37.3 75.3 69.9 43.2 38.6 - 39.5 45.2 43.4
VILA1.5-13B [lin2023vila] 69.9 - 42.5 74.9 66.3 44.3 37.9 - 48.1 50.2 53.3
CogVLM2-19B [hong2024cogvlm2] 73.4 81.0 38.6 80.5 - 60.4 44.3 - - 66.0 62.9
LLaVA-OneVision-7B [li2024llava] 81.4 80.0 63.2 80.8 - 57.5 48.8 24.1 53.0 65.4 69.9
InternVL2-8B [chen2023internvl] 83.8 83.3 58.3 81.7 81.2 54.2 49.3 29.0 50.9 67.3 64.2
MiniCPM-V2.5-8B [yao2024minicpm] 78.4 - 54.3 77.2 74.2 52.8 45.8 - - 61.4 63.5
MiniCPM-V2.6-8B [yao2024minicpm] 82.1 - 60.6 - - 60.0 49.8 27.2 55.2 65.7 65.0
Qwen2-VL-7B [wang2024qwen2vl] 77.5 83.0 58.2 83.0 80.5 62.0 54.1 30.5 53.8 68.6 68.5
InternVL2.5-8B [chen2024expanding] 84.8 84.8 64.4 84.6 82.6 62.8 56.0 34.3 54.8 69.7 70.1
Qwen2-VL-7B-GenRecal (InternVL2.5-78B) 93.9 95.3 68.8 88.4 87.4 70.4 65.6 49.6 64.3 70.7 79.2
InternVL2.5-8B-GenRecal (InternVL2.5-78B) 93.0 93.6 74.9 89.5 88.2 73.2 68.1 48.8 65.3 72.3 81.4
VILA1.5-3B [lin2023vila] 57.9 - 31.6 - - 38.8 34.2 - 39.7 41.4 53.2
Phi-3.5-Vision-4B [abdin2024phi] 77.8 81.8 43.9 76.0 66.1 43.2 43.0 19.7 58.3 62.2 53.6
InternVL2-4B [chen2023internvl] 78.9 81.5 58.6 78.6 73.9 51.0 34.3 32.7 46.1 63.9 60.7
InternVL2.5-4B [chen2023internvl] 81.4 84.0 60.5 81.1 79.3 60.6 52.3 32.7 50.8 66.9 64.3
InternVL2.5-4B-GenRecal (InternVL2.5-78B) 90.0 91.1 70.4 88.4 86.2 66.1 58.3 45.9 63.5 69.3 75.2
InternVL2-2B [chen2023internvl] 74.1 76.2 46.3 73.2 70.9 39.5 34.3 18.2 43.8 60.0 57.3
Qwen2-VL-2B [wang2024qwen2vl] 60.2 73.5 43.0 74.9 73.5 49.5 41.1 21.2 45.2 61.2 62.6
Aquila-VL-2B [chen2024expanding] 75.0 76.5 59.0 - - 43.8 47.4 - 34.1 63.0 65.0
InternVL2.5-2B [chen2024expanding] 74.9 79.2 51.3 74.7 71.9 60.8 43.6 23.7 44.0 60.0 60.1
Qwen2-VL-2B-GenRecal (InternVL2.5-78B) 89.1 90.9 60.5 82.9 81.0 57.3 52.9 37.5 53.4 65.5 72.8
InternVL2.5-2B-GenRecal (InternVL2.5-78B) 91.8 90.7 62.5 84.1 80.4 61.2 59.2 36.6 55.5 67.3 73.0
LLaVA-OneVision-0.5B [li2024llava] 57.1 61.4 34.8 61.6 55.5 32.2 31.4 - 52.1 45.7 55.6
InternVL2-1B [chen2023internvl] 64.1 72.9 37.7 65.4 60.7 32.7 36.7 14.8 38.6 54.3 50.3
InternVL2.5-1B [chen2024expanding] 69.3 75.9 43.2 70.7 66.3 48.8 40.9 19.4 42.0 59.0 57.5
InternVL2.5-1B-GenRecal (InternVL2.5-78B) 82.9 89.4 56.5 80.1 74.4 54.4 45.6 28.7 48.5 65.1 70.8
Figure 4: Overview of regularization simultaneously done with first training stage-alignment.
However, existing works are largely constrained to VLMs sharing identical token types, including vocabulary size, token segmentation, and token index ordering. When the large and small VLMs employ different tokenizers or image embedding strategies, distillation becomes problematic—KL divergence cannot be computed due to mismatched output token counts, and token indices may no longer correspond to semantically equivalent meaning. Overcoming this limitation would represent a significant advancement, broadening the applicability and robustness of distillation techniques for diverse VLM architectures.
3 GenRecal: Generation after Recalibration
3.1 Model Architecture & Various Accommodation
We briefly illustrate the overall model architecture and training procedure in Fig. 3. GenRecal consists of three main components: a large (teacher) VLM, a small (student) VLM, and a Recalibrator module. For the large VLMs, we select models with over 72B parameters to ensure competitive performance: NVLM-72B (Qwen2-72B) [nvlm2024], Qwen2-VL-72B (Qwen2-72B) [wang2024qwen2vl], InternVL2-76B (Llama3-70B) [chen2024far], and InternVL2.5-78B (Qwen2.5-72B) [chen2024expanding]. For the small VLMs, we adopt various model sizes, including Qwen2-VL-2B/7B (Qwen2-2B/7B) and InternVL2.5-1B/2B/4B/8B (InternLM2.5-1.8B/7B and Qwen2.5-0.5B/3B/72B). Note that the parentheses denote the LLMs used within each VLM. Throughout this section, we divide a VLM architecture into four key modules: the vision encoder, vision projector, VLM-body, and VLM-head (i.e. language head). Recalibrator is designed to acquire shared feature representations within a common latent space. It consists of two decoder blocks and two projection layers, as depicted in Fig. 3(c). We denote the decoder blocks as Rec-body, and the projection layers as Rec-proj-pre and Rec-proj-post. Rec-body adopts the same structural design as the student VLM’s decoder, while both Rec-proj-pre and Rec-proj-post are implemented as single linear layers.
3.2 Design and Propagation of Recalibrator
To construct a shared representation space, we first extract and align hidden features from both the large and small VLMs. Since the hidden space of a large VLM typically encompasses a richer vocabulary and more expressive representations, we define it as the shared latent space which the small VLM is projected into. However, direct alignment of the two VLMs’ tokens within this space is infeasible due to discrepancies in vocabulary size, token split, and token index ordering. To overcome this challenge, we draw inspiration from the principle of autoregressive modeling, wherein a model predicts answer tokens for the given question tokens.
Specifically, we feed the same question-answer pair into both the large and small VLMs, obtaining their VLM-body outputs as:
$=\textit{VLM-body}_{l}([q_{l},a_{l}]),\quad$ $(l:\text{large})$ (1)
$[z_{q_{s}},z_{a_{s}}]$ $=\textit{VLM-body}_{s}([q_{s},a_{s}]),\quad$ $(l:\text{small})$
where $q$ and $a$ denote the question and answer tokens produced by each model’s vision encoder, vision projector, and word embedding [mikolov2013efficient]. We then extract the question features $z_{q_{s}}$ from the small VLM and the answer features $z_{a_{l}}$ from the large VLM, and concatenate them to form a joint sequence: $z_{q_{s}},z_{a_{l}}]$, which is then passed through Recalibrator. Because directly embedding both features into a shared space is infeasible due to different token types, we instead design an autoregressive loss that predicts the answer token index of the large VLM given the question token index of the small VLM. This objective encourages Recalibrator to effectively project the small VLM’s features into the large VLM’s latent space.
Table 2: (b) Comparison of GenRecal with large-scale open-source and closed-source VLMs on challenging benchmarks: MMB [liu2023mmbench], $M$-Vet [yu2023mm], $M$-Vet-v2 [yu2024mm], MMMU [yue2023mmmu], MMMU-Pro [yue2024mmmu], MMStar [chen2024we], AI2D [kembhavi2016diagram], ChartQA [masry2022chartqa], SEED-2-Plus [li2024seed], MathVista [lu2023mathvista], BLINK [fu2024blink], RealWorldQA (RWQA).
VLMs MMB $M$-Vet $M$-Vet-v2 MMMU MMMU-Pro MMStar AI2D ChartQA SEED-2-Plus MathVista BLINK RWQA
NVLM-72B [nvlm2024] - 58.9 - 59.7 - 63.7 85.2 86.0 68.4 66.6 48.0 69.9
LLaVA-OneVision-72B [li2024llava] 85.8 60.6 - 56.8 31.0 65.8 85.6 83.7 - 67.5 55.4 71.9
Molmo-72B [deitke2024molmo] - 61.1 - 54.1 - 63.3 83.4 87.3 - 58.6 - 73.7
Qwen2-VL-72B [wang2024qwen2vl] 86.5 74.0 68.7 64.5 46.2 68.3 88.1 88.3 72.3 70.5 60.5 77.8
InternVL2-76B [chen2023internvl] 86.5 65.7 68.4 62.7 40.0 67.4 87.6 88.4 69.7 65.5 56.8 72.2
InternVL2.5-78B [li2024llava] 88.3 72.3 65.5 70.1 48.6 69.5 89.1 88.3 71.3 72.3 63.8 78.7
Claude-3.5-Sonnet [claude3series2024] 82.6 70.1 71.8 68.3 51.5 65.1 81.2 90.8 71.7 67.7 60.1 60.1
Gemini-1.5-Pro [team2023gemini] 73.9 64.0 66.9 62.2 46.9 59.1 79.1 87.2 70.8 63.9 59.1 67.5
GPT-4o (0513) [gptsyscard] 83.4 69.1 71.0 69.1 51.9 64.7 84.6 85.7 72.0 63.8 68.0 75.4
InternVL2.5-8B-GenRecal (NVLM-72B) 88.8 63.9 59.3 60.3 34.7 65.1 88.5 84.3 69.3 67.5 55.6 72.5
InternVL2.5-8B-GenRecal (InternVL2-76B) 89.0 72.4 65.0 64.6 42.8 65.8 92.1 90.4 70.9 71.9 57.7 75.5
InternVL2.5-8B-GenRecal (Qwen2-VL-72B) 89.5 71.4 62.6 65.6 43.8 66.1 92.2 87.5 71.3 71.4 61.4 78.8
InternVL2.5-8B-GenRecal (InternVL2.5-78B) 89.5 73.2 67.2 68.1 48.8 67.8 93.0 93.6 72.3 74.9 65.3 81.4
However, the hidden dimensions of large and small VLMs often differ. To address this, in Recalibrator, we first apply Rec-proj-pre to $z_{a_{l}}$, aligning its dimensionality to that of $z_{q_{s}}$. The concatenated sequence $z_{q_{s}},\textit{Rec-proj-pre}(z_{a_{l}})]$ is then propagated through Rec-body decoder blocks. The output of Rec-body is subsequently passed through Rec-proj-post to restore the hidden dimensionality of the large VLM. Finally, the answer part of the resulting features $r_{q_{s}},r_{a_{l}}]=\textit{Recalibrator}([z_{q_{s}},z_{a_{l}}])$ are fed into VLM-head of the large VLM to compute the autoregressive loss:
$\mathcal{L}_{ar}=\text{CE}(\textit{VLM-head}_{l}(r_{a_{l}}),gt_{l}),$ (2)
where $gt_{l}$ means answer token index from large VLM. Additionally, we employ a KL divergence objective (see Algo. 1 in Appendix E) to further enhance Recalibrator’s projection capability. This process enables the large VLM to interpret the hidden features of the small VLM, thus establishing a bridge for general-purpose distillation across heterogeneous token types.
Figure 5: Distillation performance on MMMU [yue2023mmmu] for various pairings os teacher and student VLM. Each cell indicates the resulting score when using the corresponding teacher (rows) and student (columns) model sizes.
3.3 Training Process
To achieve general-purpose distillation, we adopt a three-stage training process. In the first stage, only Recalibrator is trained while keeping all parameters of both the large and small VLMs frozen, as illustrated in Fig. 3(a). During this stage, we compute both the autoregressive loss and the KL divergence. This step primarily serves to align the feature representations between the large and small VLMs. The detailed procedures for computing the loss functions are presented in Algo. 1 in Appendix E, where the purple components indicate the trainable parameters.
An important finding in this stage is the effectiveness of a regularization term (see Fig. 4), which plays a crucial role in preventing the feature representations from drifting too far from those of the large VLM. Without this regularization, the final distillation performance drops significantly, as shown in the experimental section. This degradation occurs due to insufficient feature alignment between the models. Therefore, we introduce this regularization term to better harmonize the feature representations of the large and small VLMs. The detailed implementation of this regularization is provided in Algo. 2 in Appendix E. Note that the first stage jointly optimizes both objectives from Algo. 1 and 2.
In the second stage (Algo. 3 in Appendix E), we retain the same loss functions from the first stage and additionally incorporate the small VLM’s own autoregressive loss. As depicted in Fig. 3(b), this stage enables knowledge distillation from the large VLM to the small one by training the small VLM’s VLM-body to adapt to the shared feature representation and acquire richer knowledge. In the final stage, we remove both Recalibrator and the large VLM, and fine-tune the student VLM, allowing it to perform inference independently. Specifically, we train all parameters of the student VLM except for the vision encoder through supervised fine-tuning (SFT) to further enhance its instruction-following capability.
4 Experiments
Figure 6: An overview of our training pipeline, illustrating both the question prompt and the measurement/legend annotations (top), followed by t-SNE visualizations (bottom) of teacher and student VLM pairings at the initial and final training stages. The question prompt (upper-left) shows the format of the question, while the measurement and legend box (upper-right) shows key model components to measure. Each scatter plot in the lower panels corresponds to a different combination of teacher and student VLM sizes, capturing how the learned representations evolve from early to later training iterations.
4.1 Implementation Detail
To ensure successful reproducibility, we outline three key technical aspects and details of GenRecal: (a) the structure detail of Recalibrator and (b) specific details of training and evaluation.
(a) Structure Detail of Recalibrator. This is composed of two transformer decoder blocks (Rec-body) and two projectors (Rec-proj-pre and Rec-proj-post). The decoder blocks follow all configurations of small VLMs’ decoder block such as causal mask, hidden dimensions, number of heads, and FFN structure. In addition, the projectors include the single layer linear module. For propagation in Recalibrator, there are some critical techniques that should be considered: new positional embedding (NPE) should be assigned to input sequence for Recalib-body. This is because the input sequence is based on question tokens in small VLMs and answer tokens in large VLMs, so they naturally have different type of positional embeddings and different position-ids as well. Therefore, we employ another RoPE [su2024roformer] to realign their positional embeddings, and their position-ids are re-assigned as well. Lastly, we apply additional layer-norm [ba2016layernormalization] into the output features of Recalibrator for stable adaptation, before meeting VLM-head of large VLM.
[…]
4.2 Validation and Analysis of GenRecal
Higher Size, Higher Performance. Throughout its realization of GenRecal, Tab. 1 demonstrates that GenRecal outperforms not only previous standard model size VLMs but also smaller size VLMs, highlighting that GenRecal generalizes well across different model sizes. From this table, we can say that if we choose more high-performing small VLMs like InternVL2.5-8B [chen2024expanding] significantly impacts the overall distillation results as well. Furthermore, when we choose same large VLMs, selecting more smaller size VLMs makes the lower distillation performances in the end. It implies that employing more capable small VLMs is crucial for achieving higher distillation performances.
Table 3: Comparing the final distillation performances from the teacher model: InternVL2.5-78B [chen2024expanding] with and without the regularization term (denoted as “Reg”). Note that, the student models are each InternVL2.5-8B, 4B, 2B, and 1B.
VLMs Reg MMB MathVista $M$-Vet MMMU
InternVL2.5-8B - 84.6 64.4 62.8 56.0
w. GenRecal 88.2 69.8 63.5 58.9
w. GenRecal 89.5 74.9 73.2 68.1
InternVL2.5-4B - 81.1 60.5 60.6 52.3
w. GenRecal 82.9 61.7 61.1 53.6
w. GenRecal 88.4 70.4 66.1 58.3
InternVL2.5-2B - 74.7 51.3 60.8 43.6
w. GenRecal 75.8 53.2 60.9 45.4
w. GenRecal 84.1 62.5 61.2 59.2
InternVL2.5-1B - 70.7 43.2 48.8 40.9
w. GenRecal 72.6 45.9 49.8 41.1
w. GenRecal 80.1 56.5 54.4 45.6
Table 4: Comparing SFT, traditional distillations, and GenRecal where we use equal training dataset (Section 4.1(b)) and choose the teacher (Qwen2-VL-72B [wang2024qwen2vl]) that has the same token type of student.
VLMs MMB MathVista $M$-Vet MMMU
Qwen2-VL-7B 83.0 58.2 62.0 54.1
w. SFT 84.3 60.5 64.2 56.3
w. MiniLLM 84.4 61.3 65.1 57.4
w. DistiLLM 84.8 61.5 65.3 57.9
w. LLaVA-KD 85.0 61.8 65.9 58.2
w. GenRecal 87.8 69.5 67.8 64.2
Qwen2-VL-2B 74.9 43.0 49.5 41.1
w. SFT 76.3 45.8 51.8 44.3
w. MiniLLM 77.5 46.2 52.5 45.3
w. DistiLLM 77.9 46.6 53.0 46.1
w. LLaVA-KD 78.3 46.9 53.7 46.9
w. GenRecal 82.9 58.3 57.1 51.4
Figure 7: Depicting the process of (a) computing cosine similarity, (b) the result without regularization or (c) with regularization. Because the number of the embedded tokens for small and large VLMs are naturally different due to vocabulary size, token split, and token ordering, therefore we do average the output tokens and compute cosine similarity.
Next, we examine the effect of changing large VLMs to determine whether the aforementioned property holds consistently across different large VLMs. As shown in Fig. 2, we can see that the results remain consistent, indicating that selecting more powerful large VLMs leads to greater performance improvements. To further investigate it, we utilize InternVL2.5 series [chen2024expanding] and conduct experiments with various combinations of large and small VLM sizes, as illustrated in Fig. 5. The results also consistently indicates that choosing larger size teacher and student VLMs eventually leads to higher distillation performances.
[…]
Importance of Regularization. We conduct an ablation study that investigates how important the regularization term is for the final distillation performances. We first remove the regularization term and analyze the differences before and after its removal. To illustrate this in detail, we simply prepare 10 different question-answer pairs and input them through VLM-body of both small and large VLMs. Afterwards, we propagate their output features into Recalibrator, as described in Fig. 7(a). We then compute a $10\times 10$ cosine similarity matrix between Recalibrator’s output features from the small and large VLMs. To ensure reliable cosine similarity values, we use all training dataset samples and compute the average of the cosine similarity values for all the samples.
Table 5: FLOPs comparison and several ablation studies on various configurations influencing the performance of GenRecal. Note that the teacher VLM in (b), (c), (d), (e), and (f) is InternVL2.5-78B [chen2024expanding], and the student VLM in (b) is InternVL2.5-8B [chen2024expanding].
(a) Model Size and FLOPs Comparison
Small VLM Large VLM Recalibrator Size
Qwen2-VL-7B ($25.7\times 10^{12}$) Qwen2-VL-72B ($260\times 10^{12}$) 524.8M ($2.27\times 10^{12}$)
InternVL2.5-8B ($27.1\times 10^{12}$) InternVL2.5-78B ($269\times 10^{12}$) 503.3M ($2.2\times 10^{12}$)
InternVL2.5-8B ($27.1\times 10^{12}$) Qwen2-VL-72B ($260\times 10^{12}$) 503.3M ($2.2\times 10^{12}$)
Qwen2-VL-7B ($25.7\times 10^{12}$) InternVL2.5-78B ($269\times 10^{12}$) 524.8M ($2.27\times 10^{12}$)
(d) Controlling GenRecal Dataset Size
Dataset Size MMB MathVista $M$-Vet MMMU Avg
9M 89.5 74.9 73.2 68.1 76.4
7M 89.4 74.7 73.0 67.8 76.2
5M 89.3 74.5 72.8 67.6 76.0
3M 88.9 73.6 71.9 66.6 75.3
1M 86.6 68.6 67.0 60.9 70.8
0M 84.6 64.4 62.8 56.0 67.0
(g) Compatibility with LLMs
LLMs MMLU GPQA MATH Avg
Qwen2-7B [yang2024qwen2] 70.8 38.4 48.6 52.6
Qwen2-7B-GenRecal (Llama-3-70B) 78.4 38.9 49.1 55.5
Llama-3-70B [dubey2024llama] 82.5 39.5 50.4 57.5
InternLM2.5-7B [cai2024internlm2] 72.8 38.4 60.1 57.1
InternLM2.5-7B-GenRecal (Qwen2.5-72B) 83.7 43.6 61.8 63.0
Qwen2.5-72B [yang2024qwen2] 86.1 45.9 62.1 64.7
(b) Controlling Recalibrator Depth
Depths MMB MathVista $M$-Vet MMMU Avg
20 89.9 75.1 73.6 68.3 76.7
15 89.8 75.2 73.5 68.4 76.7
10 89.6 75.0 73.4 68.3 76.6
5 89.4 74.8 73.3 68.2 76.4
2 89.5 74.9 73.2 68.1 76.4
1 88.0 73.0 72.0 66.5 74.9
(e) w/ and w/o Fine-Tuning on Our Dataset
VLM Finetuned MMB MathVista $M$-Vet MMMU Avg
InternVL2.5-78B [chen2024expanding] 88.3 72.3 72.3 70.1 75.8
InternVL2.5-78B-SFT 89.4 75.2 74.2 71.7 77.6
InternVL2.5-8B [chen2024expanding] 84.6 64.4 62.8 56.0 67.0
InternVL2.5-8B-SFT 86.5 65.2 65.0 59.9 69.2
InternVL2.5-8B-GenRecal 89.5 74.9 73.2 68.1 76.4
(h) Compatibility with Different VLM Family
VLM $M$-Vet MathVista MMMU Avg
Ovis1.6-Gemma-9B [lu2024ovis] 65.0 67.3 55.0 62.4
Ovis1.6-Gemma-9B-GenRecal (InternVL2.5-78B) 74.0 75.3 67.7 72.3
InternVL2.5-8B [chen2024expanding] 62.8 64.4 56.0 61.1
InternVL2.5-8B-GenRecal (Ovis1.6-Gemma-27B) 68.5 69.9 61.2 66.5
(c) Effect of NPE
VLM NPE MMB MathVista $M$-Vet MMMU Avg
Qwen2-VL-7B-GenRecal 79.2 60.0 64.0 63.9 66.8
Qwen2-VL-7B-GenRecal 88.4 68.8 70.4 65.6 73.3
InternVL2.5-8B-GenRecal 85.2 65.1 67.3 64.4 70.5
InternVL2.5-8B-GenRecal 89.5 74.9 73.2 68.1 76.4
(f) Comparing Cross-Tokenizer Methods
VLM MathVista MMB $M$-Vet MMMU Avg
InternVL2.5-8B 64.4 84.6 62.8 56.0 67.0
InternVL2.5-8B-ULD [boizard2025towards] 66.5 86.2 64.1 60.8 69.4
InternVL2.5-8B-MOT [cui2025multi] 68.4 87.5 65.0 62.3 70.8
InternVL2.5-8B-GenRecal 74.9 89.5 73.2 68.1 76.4
(i) Compatibility with Recent VLMs
VLM MMB MathVista $M$-Vet MMMU Avg
Qwen3-VL-8B [yang2025qwen3] 86.8 77.2 74.5 69.6 77.0
Qwen3-VL-8B (Qwen3-VL-32B) 89.9 80.1 77.6 72.7 80.1
Qwen3-VL-8B (InternVL3.5-38B) 89.7 80.3 77.4 72.5 80.0
InternVL3.5-8B [wang2025internvl3] 86.5 78.4 83.1 73.4 80.4
InternVL3.5-8B (Qwen3-VL-32B) 89.6 81.5 86.0 76.3 83.4
InternVL3.5-8B (InternVL3.5-38B) 89.4 81.3 86.2 76.5 83.4
With the regularization term included, as shown in Fig. 7(c), we observe that the diagonal values in the cosine similarity matrix are significantly higher than the off-diagonal values. This implies that the features of small and large VLMs are explicitly shared through Recalibrator, aligned with the results in Fig. 6 as well. In contrast, when the regularization term is removed, as depicted in Fig. 7(b), the diagonal values are comparable with the off-diagonal values. This indicates that Recalibrator fails to explicitly align large and small VLMs’ features, preventing feature matching and sharing. As a result, this leads to a significant drop in final distillation performances, as reported in Tab. 4. Insufficient alignment of large and student VLMs reduces efficient knowledge transfer during distillation. These findings highlight the crucial role of the regularization term in feature alignment and knowledge transfer for general-purpose distillation.
4.3 Training Computation and Configurations
In Tab. 5(a), we compare the model sizes and training FLOPs of the small VLM, large VLM, and Recalibrator to assess whether Recalibrator introduces additional computational overhead during training. The results show that Recalibrator’s FLOPs are substantially lower than those of both the large and small VLMs. Note that Recalibrator is removed at inference time, thus incurring no additional computational cost during inference.
We further conduct a series of ablation studies to examine various configurations affecting the distillation performance of GenRecal. As shown in Tab. 5(b), we vary the decoder depth of Recalibrator from 1 to 20 and select two depths based on a trade-off between computational efficiency and performance. Additionally, Tab. 5(c) demonstrates the importance of employing a new positional embedding (NPE) to handle newly introduced position-ids and to integrate RoPE [su2024roformer] from both large and small VLMs, which have different token types.
To analyze the effect of dataset scale, we control the dataset size from 1M to 9M samples, as shown in Tab. 5(d). The performance improvements between 5M and 9M samples are marginal, suggesting that a 5M dataset is a practical choice for users with limited training resources. Furthermore, Tab. 5(e) shows that fine-tuning the large VLM on our dataset yields further gains, indicating that the small VLM trained under the same setting has the potential to match or even surpass the large VLM’s performance.
We also compare GenRecal with recent cross-tokenizer distillation methods in Tab. 5(f). These baselines perform feature-level distillation via cross-token matching in the logit space by using classical method of Wasserstein distance [boizard2025towards] or optimal transport [cui2025multi], whereas GenRecal explicitly introduces Recalibrator to align hidden features for semantic correspondence. As observed, GenRecal achieves a substantial improvement over prior approaches.
Finally, Tab. 5(g)–(i) demonstrate GenRecal’s strong compatibility across different LLMs, VLM family, and recently released VLMs. In all cases, GenRecal consistently delivers superior performance, validating its general applicability to diverse model architectures.
4.4 Discussion
Cross-Token Matching. We can divide cross-token matching into two cases: (a) word embedding (vocabulary) matching and (b) semantic feature matching. Fundamentally, case (a) is infeasible because although the input tokens are aligned, the backbone decoder still cannot understand different word embeddings without retraining the entire model from scratch.
For case (b), UID [boizard2025towards] and MOT [cui2025multi] have recently been proposed as solutions for cross-token matching, which we compared in Tab. 5(f). However, these approaches suffer from potential information loss since they cannot properly handle token-split indices, leading to mismatched output lengths during distillation. To compensate, they just rely on zero-padding or truncation to enforce equal token lengths. In contrast, GenRecal introduces an explicit Recalibrator that produces equal-length token outputs fed into VLM-head of large VLM, thereby preventing information loss.
Traditional Distillation. We further compare GenRecal with traditional distillation methods under identical token types for both teacher and student VLMs. Specifically, we use Qwen2-VL series [wang2024qwen2vl] as the teacher (72B) and student (7B or 2B) VLMs. Interestingly, even when limited to the same token types, Tab. 4 shows that GenRecal outperforms both SFT and traditional distillation by a substantial margin. Note that MiniLLM [gu2024minillm] employs reverse KL divergence, DistiLLM [ko2024distillm] uses a skewed variant, and LLaVA-KD [cai2024llava] applies normal KL divergence within a three-stage framework. These results suggest even when the large and small VLMs share identical token types, explicit feature alignment remains crucial to reduce representational disparity.
Moreover, when constructing GenRecal, VLM-head of the large VLM is exclusively used for distillation into the small VLM as illustrated in Fig. 3(a) and (b), whereas traditional distillation methods employ the student’s VLM-head to absorb the teacher’s knowledge. This design enables GenRecal to capture richer and more expressive information, as the large VLM’s VLM-head inherently possesses a higher hidden dimensionality than that of the small VLM.
5 Conclusion
We present a new, efficient family of vision-language models (VLMs) called Generation after Recalibration (GenRecal). At its core, Recalibrator enables token types-compatible general-purpose distillation. For future improvements, we will expand intermediate layer’s Recalibrator to embed fine-grained and sequential knowledge, or multiple VLMs-source distillation. We envision GenRecal as a pivotal framework in the field of distillation, providing a versatile scheme.
Appendix A Token Type Examples
Appendix B Possible Distillation Pair for VLM Combinations: Traditional Distillation versus GenRecal
Figure 8: We explore the range of distillation combinations between teacher and student VLMs using two approaches: (a) traditional distillation [cai2024llava] and (b) our proposed model, GenRecal. Unlike traditional distillation—which support only a limited set of pairings—GenRecal offers the flexibility to select any model for distillation, thereby enabling a more versatile and comprehensive distillation framework.
Appendix C Extended Related Works
Evolution of Vision-Language Models.
Since the emergence of visual instruction tuning [liu2023visual], there surge many variations of VLMs such as LLaVA-1.5 [liu2023improved], InstructBLIP [dai2023instructblip], ShareGPT-4V [chen2023sharegpt4v], MiniGPT-v2 [chen2023minigpt] [liu2023improved, dai2023instructblip, chen2023sharegpt4v, chen2023minigpt, zhu2023minigpt], and they initially start with a standard model size of 7B. After that period, numerous VLMs [liu2024llavanext, mckinzie2024mm1, li2024mini, dong2024internlm4khd] have divided an image into multiple sub-regions to focus on the image’s details so that visual perception is enhanced. CoLLaVO [lee2024collavo] and MoAI [lee2024moai] utilize computer vision models directly for visual capability, and Mini-Gemini [li2024mini], MoVA [zong2024mova], and Eagle [shi2024eagle] employ multiple vision encoders such as CLIP [clip], ConvNext [woo2023convnext], DINO-v2 [oquab2023dinov2], and SAM [kirillov2023segment]. In parallel, Meteor [lee2024meteor] explores the efficient way of learning complex reasoning abilities, and TroL [lee2024trol] and Phantom [lee2024phantom] investigate propagation modification for how we can embed vision-language knowledge as much as possible despite using the same architectures. More recently, many VLMs like Molmo-72B [deitke2024molmo], LLaVA-OneVision-72B [li2024llava], NVLM-72B [nvlm2024], Qwen2-VL-72B [wang2024qwen2vl] and InternVL2.5-78B [chen2024expanding] have employed large-scale language models such as Qwen2/Qwen2.5-72B [yang2024qwen2]. Thanks to scaling laws [kaplan2020scaling, chung2022scaling], they have closely reached the performances of GPT-4V [gptsyscard] and Claude-3.5 Sonnet [claude3series2024]. Therefore, developing VLMs with large-scale language models is getting to a standard these days.
Knowledge Distillation.
Knowledge distillation [hinton2015distilling] provides an effective framework for transferring the rich representations of large, high-performing models to smaller, efficient counterparts while preserving performance. Early studies primarily focused on aligning the output logits of a “teacher” model with those of a “student” [sanh2019distilbert, turc2019well]. Subsequently, the FitNet framework [romero2014fitnets] extended this idea by supervising the student using intermediate feature representations rather than final outputs. In this setup, convolutional layers project the student’s features to resemble those of the teacher, and discrepancies are minimized using an L2 loss.
Following this feature-based paradigm, later methods explored more sophisticated ways of leveraging internal representations for effective knowledge transfer. For instance, [sun2019patient, chen2021distilling, wang2024crosskd, ben2022s] demonstrated that incorporating multiple intermediate layers from the teacher significantly enhances distillation effectiveness. Probabilistic approaches such as PKT [passalis2020probabilistic] reformulated the teacher’s knowledge as a probability distribution, aligning it with the student’s outputs via KL divergence. Similarly, RKD [park2019relational] captured inter-sample relationships, while CRD [tian2019contrastive] combined contrastive learning with traditional distillation.
Multi-stage and relational strategies further broadened the field. AT [zagoruyko2016paying] utilized attention maps from several teacher layers, and FSP [yim2017gift] employed matrices derived from feature maps to guide the student. Later, SP [tung2019similarity] evaluated similarity among input samples, and OFD [heo2019comprehensive] introduced a margin-based distance metric derived from ReLU activations to capture essential information.
As the methodology matured, two main paradigms emerged. On-policy approaches [austin2021program, ko2024distillm, agarwal2024policy, gu2024minillm] dynamically sample data during training, whereas off-policy methods [ko2024distillm, sun2019patient, turc2019well, zhang2024dual, boizard2024towards] rely on pre-collected datasets. While early research mostly adopted single-teacher frameworks [agarwal2024policy, gu2024minillm, shu2024llava, muralidharan2024compact], multi-teacher distillation has gained increasing attention despite challenges such as architectural differences, vocabulary mismatches, and task misalignment [timiryasov2023baby, lee2023ensemble, wan2024fusechat]. Representative examples include DKMF [wang2021distilling] and FNKD [xu2020feature], which utilize teacher feature representations, while DGKD [son2021densely] aggregates knowledge from multiple teachers to boost student performance.
Recent work has also emphasized refining loss functions used in distillation. MiniLLM [gu2024minillm] and DistiLLM [ko2024distillm] propose modified KL divergences—such as reverse or skewed variants—to reduce overfitting, particularly for long-tail predictions. Similarly, [wu2024rethinking] introduced a dynamic loss-balancing scheme that adaptively combines conventional and reverse KL divergences during training. Finally, to enhance reasoning capabilities, recent studies integrate chain-of-thought (CoT) supervision. Methods such as [hsieh2023distilling, tian2024beyond] leverage detailed reasoning traces from larger models, while TinyLLM [tian2024beyond] aggregates multi-teacher reasoning to enrich training signals and improve generalization.
Appendix D Loss Graphs for Recalibrator
[…]
Appendix E Implementation Details of GenRecal for Algorithms
Algorithm 1 First stage Loss Functions
1:Input: ($q_{l}$, $a_{l}$), ($q_{s}$, $a_{s}$), $gt_{l}$ (label index)
[…]
7:Return: $\mathcal{L}_{ar}+\mathcal{L}_{kl}$
Algorithm 2 First stage Regularization
1:Input: ($q_{l}$, $a_{l}$), ($q_{s}$, $a_{s}$), $gt_{l}$ (label index)
[…]
6:Return: $\mathcal{L}_{ar}+\mathcal{L}_{kl}$
Algorithm 3 Second stage Loss Functions
1:Input: ($q_{l}$, $a_{l}$), ($q_{s}$, $a_{s}$), ($gt_{l}$, $gt_{s}$) (label index)
[…]
8:Return: $\mathcal{L}_{ar}+\mathcal{L}_{kl}$
Recalibrator’s propagation rule slightly changes depending on the hidden dimension needed. Its propagation of Fig. 3(c) and 4 correspond to that of Algorithm 1 and Algorithm 2, respectively. In Figure 3(c), Recalibrator takes small VLM’s tokens for the question and large VLM’s tokens for the answer. In contrast, in Figure 4, only large VLM’s tokens (question and answer) are fed into Recalibrator. Rec-body has the hidden dimension of the small VLM, so we should match this hidden dimension of large VLM’s tokens to that of small VLM’s ones. This is where Rec-proj-pre comes into play. This ensures compatibility with Rec-body’s hidden dimension.
Appendix F Dataset Composition and Analysis for Building GenRecal
We gather 9M visual instruction tuning dataset covering a wide range of vision-language capabilities such as general visual question answering, dense image captioning, chart/diagram/document understanding, common-sense knowledge, science and math understanding, and multi-dimensional reasoning. Our dataset includes LLaVA-OneVision [li2024llava], MMC [liu2023mmc], DenseFusion [li2024densefusion], Cambrian [tong2024cambrian], GPT-4V-filtered synthetic data of SA-1B [kirillov2023segment] and Infinity-MM [gu2024infinity], Finance-QA [Sujet-Finance-QA-Vision-100k], Wikipedia knowledge [li2024cvlm], InfoSeek [chen2023can], science & mathematical reasoning (SMR) [zhang2024beyond], document-downstream/reasoning [hu2024mplug], WildVision [lu2024wildvision], SROIE [huang2019icdar2019], RLAI-F [yu2024rlaif], M3CoT [chen-etal-2024-m3cot], LLaVAR [zhang2023llavar], KonIQ [hosu2020koniq], iNaturalist2018 [van2018inaturalist]. For further analysis, we particularly categorize the visual instruction tuning dataset that we use to build GenRecal into three domains: ‘Knowledge’, ‘Science & Math’, and ‘Chart & Document’. Based on these categories, we conduct dataset composition analysis by removing each of the three domains. This implies that MMMU [yue2023mmmu] requires VLMs’ ability for the domain of ‘Knowledge’ more than the others, while MathVista [lu2023mathvista] requires their capabilities for ‘Science & Math’.
Figure 10: Accuracy trends across different datasets: (Left) MMMU [yue2023mmmu], and (Right) MathVista [lu2023mathvista] over three training stages. The performance is shown for different category of visual instruction tuning dataset: “All” (without any exclusion), “w.o. Knowledge”, “w.o. Science & Math”, and “w.o. Chart & Document”. The results indicate the impact of these exclusions on accuracy progression, highlighting how the absence of specific knowledge domains affects. Note that, Note that each data point within the same stage represents 20% of the training progress, with five data points measured per stage.
  • Knowledge: GPT-4V filtered synthetic of SA-1B [kirillov2023segment] and Infinity-MM [gu2024infinity], Wikipedia Knowledge [li2024cvlm], Wild-Vision, InfoSeek [chen2023can], WildVision [lu2024wildvision], and iNaturalist2018 [van2018inaturalist]
  • Science & Math: LLaVA-OneVision (subset) [li2024llava], Cambrian (subset) [tong2024cambrian], Finance-QA [Sujet-Finance-QA-Vision-100k], SMR [zhang2024beyond], M3CoT [chen-etal-2024-m3cot], and KonIQ [hosu2020koniq]
  • Chart & Document: LLaVA-OneVision (subset) [li2024llava], Cambrian (subset) [tong2024cambrian], MMC [liu2023mmc], RLAI-F [yu2024rlaif], document-downstream/reasoning [hu2024mplug], SROIE [huang2019icdar2019], and LLAVAR [zhang2023llavar]